OpenStudy (anonymous):

What are formed by perpendicular lines? - 90 angles or - 180 angles

OpenStudy (saifoo.khan):

What do you see?

Parth (parthkohli):

- 90 angles.

OpenStudy (precal):

look at the slopes you are looking for opposite signs (one positive and the other negative) and for reciprocals. example a and 1/a y=2x+3 y=(-1/2)x-7 are perpendicular m=2 m=-1/2